# The equation of Binet in classical and relativistic orbital mechanics

### v1

@@ Appendix B Norm of the tangent vector of a geodesic
As a result of this, the squared norm of the tangent vector $v^{\mu}$ is constant along the geodesic. If $\lambda=\tau$, this constant equals $c^{2}$ for massive particles and 0 for photons, reflecting the definition of proper time and the fact that photons lie on the boundary of spacetime’s causal structure.
@@
where $C$ is a constant and where (38) has been used. For the case $\lambda=\tau$ and $\alpha=\phi$, we have $C=-c^{2}$, and for photons, $C=0$. The pseudo-normalization condition (45) could have been anticipated from $w^{\mu}=v^{\mu}\left(d\lambda/d\phi\right)$ and the analogous expression for $w_{\mu}$. However, the intention here was to show how it arises from the non-affine geodesic equation.
@@ Appendix D Conservation laws from the geodesic equations
There are various ways to derive the conservation laws in relativistic mechanics. One elegant approach is to use the Killing vectors of the spacetime metric [carroll:2019, pp. 219 ff.]. An equivalent method is to use the geodesic equations for the coordinates that do not appear explicitly in the metric components. In the Schwarzschild-(anti-)de Sitter metric given in (26), these coordinates are $t$ and $\phi$. We begin by obtaining the first integrals of the geodesic equations, using $\tau$ as the affine parameter for these two coordinates. By substituting the values of the Christoffel symbols given in Appendix C, it is straightforward to write equations (23) for $\mu=t$ and $\mu=\phi$ with $\lambda=\tau$ as
@@
The conservation law for the angular momentum allows us to compute the pseudo-normalization condition given in equation (45) with $\lambda=\tau$ and $\alpha=\phi$. This is the only relationship involving the proper time needed in the present derivation of Binet’s equation, a derivation that uses the non-affine geodesic equations and the pseudo-normalization condition given in (45) with $\alpha=\phi$ and $\lambda=\tau$.
@@
The combination of the pseudo-normalization condition (45) with $\alpha=\phi$ and $\lambda=\tau$ with equation (53) yields the square of the derivative of $u$ with respect to $\phi$
@@ Appendix F Solution to the relativistic lightlike Binet’s equation
so that the photon orbit to be given by
@@
The phase constant $\phi_{\mathrm{in}}$ is fixed by the initial condition $r(\phi_{0})=r_{0}$ via inversion of the $\wp$-function. It is a non-trivial elliptic function and its full shape can be found in [hackmannthesis:2010].

### v2

@@ Appendix B Norm of the tangent vector of a geodesic
As a result of this, the squared norm of the tangent vector $v^{\mu}$ is constant along the geodesic. This constant equals $c^{2}$ for massive particles and 0 for photons, reflecting the definition of proper time and the fact that photons lie on the boundary of spacetime’s causal structure.
@@
where $C$ is a constant and where (38) has been used. For the case $\lambda=\tau$ (or simply $\lambda$ for photons) and $\alpha=\phi$, we have $C=-c^{2}$ (massive particles) or $C=0$ (photons). The pseudo-normalization condition (45) could have been anticipated from $w^{\mu}=v^{\mu}\left(d\lambda/d\phi\right)$ and the analogous expression for $w_{\mu}$. However, the intention here was to show how it arises from the non-affine geodesic equation.
@@ Appendix D Conservation laws from the geodesic equations
There are various ways to derive the conservation laws in relativistic mechanics. One elegant approach is to use the Killing vectors of the spacetime metric [carroll:2019, pp. 219 ff.]. An equivalent method is to use the geodesic equations for the coordinates that do not appear explicitly in the metric components. In the Schwarzschild-(anti-)de Sitter metric given in (26), these coordinates are $t$ and $\phi$. We begin by obtaining the first integrals of the geodesic equations, using $\tau$ as the affine parameter for these two coordinates (or $\lambda$ in the case of photons). By substituting the values of the Christoffel symbols given in Appendix C, it is straightforward to write equations (23) for $\mu=t$ and $\mu=\phi$ with $\lambda=\tau$ (or simply $\lambda$ for photons) as
@@
The conservation law for the angular momentum allows us to compute the pseudo-normalization condition given in equation (45) with $\lambda=\tau$ (or $\lambda$ for photons) and $\alpha=\phi$. This conservation law is the only relationship involving the proper time needed in the present derivation of Binet’s equation.
@@
The combination of the pseudo-normalization condition (45) with $\alpha=\phi$ and $\lambda=\tau$ (or $\lambda$ for photons) with equation (53) yields the square of the derivative of $u$ with respect to $\phi$
@@ Appendix F Solution to the relativistic lightlike Binet’s equation
so that the photon orbit is given by
@@
The phase constant $\phi_{\mathrm{in}}$ is fixed by the initial conditions. It is a non-trivial elliptic function and its full shape can be found in [hackmannthesis:2010].